### v1

@@ 6 Conclusions
Appendix A Author Contributions
The authors’ contributions are: Xiaochen Ma: codebase design, the coding leader, and manuscript writing. Xuekang Zhu: implement ObjectFormer, backobone models, and extractor modules; and manuscript writing. Lei Su: implement MVSS-Net, NCL-IML, ManTra-Net, and cleaned NIST16; and manuscript writing. Bo Du: implement PSCC-Net, Trufor, and manuscript writing. Zhuohang Jiang: implement GPU-accelerated evaluators, and manuscript writing. Bingkui Tong: implement Cat-Net, Grad-CAM, and manuscript writing. Zeyu Lei: implement ManTra-Net, SPAN, and manuscript writing. Xinyu Yang: dataset debiasing, metrics analyzing, and manuscript writing. Jiancheng Lv: General project advising. Chi-Man Pun: Project advising. Jizhe Zhou: Project supervising and manuscript writing.
Appendix B Appendix
B.1 Differences in Pretraining, Training, Evaluation, Metrics between SoTA models
@@
Table 8: The information about the pretraining and training datasets for the models we reproduced.IN, SD, PI, CO, CA2, DT, FR, Cov, IMD, CO+, RA+, NI16, Pa+, and Col correspond to ImageNet, Synthetic Dataset, pristine images, COCO dataset, CASIAv2, DEFACTO, FantasticReality, Coverage, IMD2020, COCO+ (a tampered dataset generated from COCO), RAISE+ (a tampered dataset generated from RAISE), NIST16, Paris+ (a tampered dataset generated from Paris), and Columbia, respectively.
@@
Table 9: Here is the validation dataset test metric information for the models we have reproduced. NI16, Col, CA, PSB, COV, DT, Car, CMF, IMD, DSO, VP, OF, CoG respectively represent NIST16, Columbia, CASIA, PhotoShop-battle, COVERAGE, DEFACTO, Carvalho, CoMoFoD, DSO-1, VIPP, OpenForensics, and CocoGlide.
B.2 SoTA models in model zoo
@@
ManTra-Net [41] ManTra-Net-pretrain aligns ManTraNet-pytorchfootnotemark: with our benchmark. We denormalize the input images and transfer the weight into our form, then test on five datasets.
MVSS-Net [4] MVSS-Net uses BayarConv and SRM operator as feature extractors, with ResNet50 as its backbone network. During replication, we used the code from the repository mentioned in the main text, making no major modifications. We did not use AMP during training. The model was trained for 200 epochs with an initial learning rate of 1e-4. MVSS-Net performs label prediction.
CAT-Net [21] The CAT-Net model first extracts features through convolution from both the RGB Stream and the DCT Stream, then fuses the information from these two parts. During our replication process, we primarily used the official implementation of the model structure. However, we reimplement artifact learning module in PyTorch to leverage GPU acceleration. This model does not include a label prediction function. The total number of training epochs was set to 200, with a learning rate of 0.0005.
@@
Objectformer [39] ObjectFormer uses the transformer block from the pre-trained ViT-B/16 model as the backbone. The input image size is 224. Data preprocessing involves using the FFT package to extract high-frequency information. The image and high-frequency data are passed to their respective patch embed layers and concatenated along the sequence length dimension. ObjectFormer initializes the first eight layers of the same pre-trained transformer blocks twice for the encoder and decoder parts. In the encoder, a trainable parameter serves as the query, with the concatenated features as the key and value. The encoder’s output functions as the query, key, and value for the decoder. The feature then undergoes a BCIM feature change, completing the encoder and decoder forward propagation process. After repeating this process eight times, deconvolution and linear interpolation are applied, and the output is reshaped into the mask required for prediction during training. ObjectFormer uses a cosine learning rate schedule starting from 1e-5 to 5-7, training the model for 1100 epochs on the Protocol-MVSS dataset and 2000 epochs on the Protocol-CAT dataset.
B.3 Detailed Introduction of Evaluators
@@
In the field of Image Manipulation Detection & Localization (IMDL), the standardization of evaluation metrics is essential for ensuring the comparability and reproducibility of research outcomes. The domain has currently defined seven core evaluation metrics, categorized into two major classes: image-level and pixel-level. At the image level, the evaluation metrics include the Area Under the Curve (AUC) score, F1 score, and accuracy. The pixel-level evaluation metrics further expand to include AUC, F1 score, accuracy, and Intersection over Union (IoU).
@@
However, due to the variability in dataset preprocessing methods in previous research, there is a certain degree of disorder in the existing evaluation system. For instance, differences in image size adjustment (resize) and padding strategies among models can lead to discardable content in the output prediction mask. Additionally, the inconsistency in how manipulated areas are annotated across different models—some marking manipulated areas as 1 while others as 0 has resulted in inconsistencies in the calculation methods and results of the evaluation metrics, and even completely opposite outcomes in some cases. This inconsistency greatly complicates the establishment of a unified evaluation standard and benchmark in the IMDL field.
@@
B.3.1 Confusion Matrix Acceleration
@@
B.3.2 Pixel-Level Evaluators Acceleration
@@
B.3.3 Image-Level Evaluators Acceleration
@@ B.4 GradCAM for Analysis
B.5 Additional Analysis and Insights
B.5.1 Modules Experiments Setting
Our experiments utilized the CASIAv2 dataset for training purposes, while evaluation was conducted on four different datasets to assess the generalization capabilities: CASIAv1, Columbia, NIST16, and Coverage. In addition to ViT-B/16, our experiments modified the input layer to include the feature maps extracted by the feature extractors. For instance, the DCT feature extractor’s output was three channels, which were concatenated with the original image to create a 6-channel input for the input layer. If the backbone output feature shape is inconsistent with the mask size, deconvolution was employed to ensure that the dimensions of the predicted mask matched the expected output. All models were initialized with pre-trained weights.
@@ B.5.1 Modules Experiments Setting
B.5.2 Modules Detailed Evaluation
@@
Feature Etractor None Bayar DCT FFT Sobel SRM None Bayar DCT FFT Sobel SRM
@@
Feature Etractor None Bayar DCT FFT Sobel SRM None Bayar DCT FFT Sobel SRM
@@
Average F1 0.515 0.160 0.404 0.404 0.163 0.397 0.586 0.234 0.509 0.487 0.247 0.431
Best Epoch 150 180 200 169 77 151 77 177 137 157 173 165
@@
Feature Etractor None Bayar DCT FFT Sobel SRM None Bayar DCT FFT Sobel SRM
CASIA v1 0.5169 0.0698 0.0831 0.0823 0.0662 0.0791 0.4195 0.2338 0.2989 0.2949 0.2031 0.4308
Columbia 0.3609 0.0854 0.0757 0.0969 0.0753 0.1131 0.189 0.1441 0.1275 0.2656 0.1747 0.2063
NIST16 0.2344 0.1104 0.1431 0.1225 0.1042 0.1327 0.2047 0.1415 0.156 0.163 0.1273 0.1832
Coverage 0.1365 0.0463 0.0745 0.0532 0.0551 0.063 0.0932 0.0585 0.06395 0.1028 0.0614 0.0824
Average F1 0.312 0.078 0.094 0.089 0.075 0.097 0.227 0.144 0.162 0.207 0.142 0.226
@@
B.5.3 Details of the NIST16 Dataset Cleansing Process
Dataset Type Manipulation type
auth tamp cope-move splice remove
CASIA v2 7491 5063 3235 1828 0
CASIA v1 0 920 459 461 0
IMD2020 414 2010 $\times$ $\times$ $\times$
NIST16 0 564 236 225 103
COVERAGE 100 100 100 0 0
Columbia 183 180 0 180 0
Fantastic Reality 16592 19423 $\times$ $\times$ $\times$
tampCOCO 0 800000 $\times$ $\times$ $\times$
compRAISE 24462 0 0 0 0
@@ B.5.4 Controversial F1 Metrics
Such a formula is controversial as it does not accurately reflect the model’s detection capabilities. In Figure 4, the model’s prediction results are completely incorrect, yet its F1 score is relatively high.
@@
In the Python sklearn library, the f1 score function has five options for calculation methods: ‘micro’, ‘macro’, ‘samples’, ‘weighted’, ‘binary’. The sample parameter is mainly used for multi-label classification problems and unsuitable for our tampering detection task.
@@ B.5.5 AUC Metrics
The curve in the AUC metric usually refers to the Receiver Operating Characteristic (ROC) curve, which shows the model’s performance in terms of correctly classified positive examples and incorrectly classified negative examples. Here, we briefly describe the calculation method of AUC, involving True Negatives (TN), True Positives (TP), False Positives (FP), and False Negatives (FN):
$AUC\propto\frac{TP}{TP+FN}*\left(1-\frac{FP}{FP+TN}\right)$ (2)
@@
The manipulated regions in an image are usually very small. Therefore, in manipulated images, the number of manipulated pixels and the number of authentic pixels are extremely unbalanced. Due to such a large skew in class distribution, although AUC performs well on many unbalanced classification problems, it still tends to provide an over-optimistic estimate of IML model performance. Moreover, AUC averages over all possible thresholds, but in real-life scenarios, we can only apply a fixed threshold. In other words, a high AUC value indicates that if we can find the optimal threshold, the model will perform well on IML tasks; however, it does not guarantee that the model has correctly found the optimal threshold. State-of-the-art models with high AUC values may perform poorly in tampering localization. Recent studies [4], [46] have also noticed similar issues with finding the optimal threshold, but they did not provide a profound solution. In this paper, we speculate that this disastrous performance is due to overfitting caused by large-scale pre-training, and the AUC metric is not sufficient to address this overfitting problem. Therefore, considering the overly optimistic estimates of the AUC metric in IML, we advocate using only the F1 score to evaluate IML models.

### v2

@@ 6 Conclusions
7 Author Contributions
The authors’ contributions are: Xiaochen Ma: codebase design, the coding leader, and manuscript writing. Xuekang Zhu: implements ObjectFormer, backobone models, and extractor modules; and manuscript writing. Lei Su: implements MVSS-Net, NCL-IML, ManTra-Net, and cleaned NIST16; and manuscript writing. Bo Du: implements PSCC-Net, Trufor, and manuscript writing. Zhuohang Jiang: implements GPU-accelerated evaluators, and manuscript writing. Bingkui Tong: implements CAT-Net, Grad-CAM, and manuscript writing. Zeyu Lei: implements ManTra-Net, SPAN, and manuscript writing. Xinyu Yang: dataset debiasing, metrics analyzing, and manuscript writing. Jiancheng Lv: general project advising. Chi-Man Pun: project advising. Jizhe Zhou: project supervisor and manuscript writing.
8 Acknowledgement
This research was supported by the Sichuan Provincial Natural Science Foundation (Grant No.2024YFHZ0355), the Fundamental Research Funds for the Central Universities (Grant No.2022SCU12072 and No.YJ2021159), and the Science and Technology Development Fund, Macau SAR (Grant 0141/2023/RIA2 and 0193/2023/RIA3). The authors would like to give special thanks to Kaiwen Feng for his attentive work in analyzing the macro-F1 issues and fixing bugs on the IMDL-BenCo codebase and Dr. Wentao Feng for the workplace, computation power, and physical infrastructure support.
Appendix A Appendix
A.1 Differences in Pretraining, Training, Evaluation, Metrics between SoTA models
@@
Table 8: The information about the pretraining and training datasets for the models we reproduced. IN, SD1, TF, CO, DID, KCMI, SD2, CA2, DT, FR, Cov, IMD, CO+, RA+, NI16, Pa+, and Col correspond to ImageNet, Synthetic Dataset 1, Trufor dataset for noiseprint++, COCO dataset, Dresden Image Database, Kaggle Camera Model Identification, Synthetic Dataset 2, CASIAv2, DEFACTO, FantasticReality, Coverage, IMD2020, COCO+ (a tampered dataset generated from COCO), RAISE+ (a tampered dataset generated from RAISE), NIST16, Paris+ (a tampered dataset generated from Paris), and Columbia, respectively.
@@
Table 9: Here is the validation dataset test metric information for the models we have reproduced. NI16, Col, CA, PSB, COV, DT, Car, GRIP, CMF, IMD, DSO, VP, OF, CoG respectively represent NIST16, Columbia, CASIA, PhotoShop-battle, COVERAGE, DEFACTO, Carvalho, GRIP Dataset, CoMoFoD, IMD2020, DSO-1, VIPP, OpenForensics, and CocoGlide.
A.2 Datasets in this paper
We follow MVSS-Protocol and CAT-Protocol to evaluate the performance of all models. Thus, eight publicly available datasets are included in this benchmark. The manipulation type, the number of manipulated images, and the number of authentic images are shown in Table 10.
Dataset Type Manipulation type
auth tamp cope-move splice remove
CASIA v2 7491 5123 3274 1849 0
CASIA v1 0 920 459 461 0
IMD2020 414 2010 $\times$ $\times$ $\times$
NIST16 0 564 236 225 103
COVERAGE 100 100 100 0 0
Columbia 183 180 0 180 0
Fantastic Reality 16592 19423 $\times$ $\times$ $\times$
tampCOCO 0 800000 $\times$ $\times$ $\times$
compRAISE 24462 0 0 0 0
A.3 SoTA models in model zoo
@@
ManTra-Net [43] ManTra-Net-pretrain aligns ManTraNet-pytorchfootnotemark: with our benchmark. We denormalize the input images and transfer the weight into our form (Since the normalization in the original repository differs), then test on five datasets.
MVSS-Net [4] MVSS-Net uses BayarConv and SRM operator as feature extractors, with ResNet50 as its backbone network. During replication, we used the code from the repository mentioned in the main text, making no major modifications. We did not use AMP (Automatic Mixed Precision) during training since it may cause NAN issues. The model was trained for 200 epochs with an initial learning rate of 1e-4. MVSS-Net performs label prediction.
CAT-Net [22] The CAT-Net model first extracts features through convolution from both the RGB Stream and the DCT Stream, then fuses the information from these two parts. During our replication process, we primarily used the official implementation of the model structure. However, we reimplement the artifact learning module in PyTorch to leverage GPU acceleration. This model does not include a label prediction function. The total number of training epochs was set to 200, with an initial learning rate of 0.0005.
@@
Objectformer [40] ObjectFormer uses the transformer block from the pre-trained ViT-B/16 model as the backbone. The input image size is 224. Data preprocessing involves using the FFT package to extract high-frequency information. The image and high-frequency data are passed to their respective patch embed layers and concatenated along the sequence length dimension. ObjectFormer initializes the first eight layers of the same pre-trained transformer blocks twice for the encoder and decoder parts. In the encoder, a trainable parameter serves as the query, with the concatenated features as the key and value. The encoder’s output functions as the query, key, and value for the decoder. The feature then undergoes a BCIM feature change, completing the encoder and decoder forward propagation process. After repeating this process eight times, deconvolution and linear interpolation are applied, and the output is reshaped into the mask required for prediction during training. Since the model released in the official repository differs from the structure mentioned in the original paper, we manually re-implement the ObjectFormer following the description from the paper. ObjectFormer uses a cosine learning rate schedule starting from 1e-5 to 5-7, training the model for 1100 epochs on the Protocol-MVSS dataset and 2000 epochs on the Protocol-CAT dataset.
A.4 More Pixel-level Performance Results
In addition to the F1 score, we also evaluated the SoTA models on pixel-level AUC and IoU, with the results presented in Table 11 and Table 12, respectively. As mentioned earlier, AUC often reflects an overoptimistic metric and struggles to accurately measure the localization precision of the model. On the other hand, IoU fails to handle the extreme imbalance between negative and positive samples, which is a common issue in IMDL datasets, where the tampered area in an image is typically quite small. Therefore, we believe that the F1 score better reflects the model’s localization performance.
Table 11: Pixel-level AUC.
Protocol-MVSS Mantra-Net[43] $\times$ 256×256 0.778 0.804 0.769 0.788 0.775 0.783
MVSS-Net[4] 512×512 0.720 0.732 0.737 0.861 0.755 0.761
CAT-Net[22] $\times$ 512×512 0.759 0.800 0.787 0.910 0.775 0.806
ObjectFormer[40] $\times$ 224×224 0.739 0.528 0.722 0.876 0.680 0.709
PSCC-Net[25] 256×256 0.697 0.814 0.725 0.833 0.778 0.769
Trufor[13] 512×512 0.911 0.928 0.820 0.946 0.840 0.889
IML-ViT[27] $\times$ 1024×1024 0.871 0.898 0.789 0.940 0.814 0.862
Protocol-CAT MVSS-Net[4] 512×512 0.870 0.933 0.790 0.912 $\times$ 0.876
CAT-Net[22] $\times$ 512×512 0.917 0.946 0.822 0.980 $\times$ 0.916
ObjectFormer[40] $\times$ 224×224 0.747 0.858 0.775 0.884 $\times$ 0.816
PSCC-Net[25] 256×256 0.884 0.946 0.828 0.919 $\times$ 0.894
Trufor[13] 512×512 0.942 0.899 0.878 0.974 $\times$ 0.924
IML-ViT[27] $\times$ 1024×1024 0.942 0.955 0.893 0.976 $\times$ 0.942
Table 12: Pixel-level IoU.
Protocol-MVSS Mantra-Net[43] $\times$ 256×256 0.066 0.348 0.135 0.092 0.062 0.141
MVSS-Net[4] 512×512 0.192 0.290 0.176 0.440 0.201 0.260
CAT-Net[22] $\times$ 512×512 0.204 0.469 0.196 0.509 0.198 0.315
ObjectFormer[40] $\times$ 224×224 0.181 0.224 0.111 0.320 0.106 0.478
PSCC-Net[25] 256×256 0.146 0.477 0.150 0.310 0.157 0.248
Trufor[13] 512×512 0.352 0.832 0.271 0.666 0.267 0.478
IML-ViT[27] $\times$ 1024×1024 0.372 0.685 0.250 0.648 0.264 0.444
Protocol-CAT MVSS-Net[4] 512×512 0.389 0.672 0.259 0.491 $\times$ 0.453
CAT-Net[22] $\times$ 512×512 0.387 0.895 0.212 0.748 $\times$ 0.561
ObjectFormer[40] $\times$ 224×224 0.188 0.670 0.203 0.320 $\times$ 0.345
PSCC-Net[25] 256×256 0.301 0.814 0.297 0.500 $\times$ 0.478
Trufor[13] 512×512 0.415 0.859 0.296 0.775 $\times$ 0.586
IML-ViT[27] $\times$ 1024×1024 0.590 0.933 0.440 0.749 $\times$ 0.678
A.5 Detailed Introduction of Evaluators
@@
In the field of Image Manipulation Detection & Localization (IMDL), the standardization of evaluation metrics is essential for ensuring the comparability and reproducibility of research outcomes. The domain currently defines seven core evaluation metrics, categorized into two major classes: image-level and pixel-level. At the image level, the evaluation metrics include the Area Under the Curve (AUC) score, F1 score, and accuracy. The pixel-level evaluation metrics further expand to include AUC, F1 score, accuracy, and Intersection over Union (IoU).
@@
However, due to the variability in dataset preprocessing methods in previous research, there is a certain degree of disorder in the existing evaluation system. For instance, differences in image size adjustment (resize) and padding strategies among models can lead to discardable content in the output prediction mask. Additionally, the inconsistency in how manipulated areas are annotated across different models—some marking manipulated areas as 1 while others as 0 has resulted in inconsistencies in the calculation methods and results of the evaluation metrics, and even wholly opposite outcomes in some cases. This inconsistency greatly complicates the establishment of a unified evaluation standard and benchmark in the IMDL field.
@@
A.5.1 Confusion Matrix Acceleration
@@
A.5.2 Pixel-Level Evaluators Acceleration
@@
A.5.3 Image-Level Evaluators Acceleration
@@ A.6 GradCAM for Analysis
A.7 Additional Analysis and Insights
A.7.1 Modules Experiments Setting
Our experiments utilized the CASIAv2 dataset for training purposes, while evaluation was conducted on four different datasets to assess the generalization capabilities: CASIAv1, Columbia, NIST16, Coverage, and IMD2020. In addition to ViT-B/16, our experiments modified the input layer to include the feature maps extracted by the feature extractors. For instance, the DCT feature extractor’s output was three channels, which were concatenated with the original image to create a 6-channel input for the input layer. If the backbone output feature shape is inconsistent with the mask size, deconvolution was employed to ensure that the dimensions of the predicted mask matched the expected output. All models were initialized with pre-trained weights.
@@ A.7.1 Modules Experiments Setting
A.7.2 Modules Detailed Evaluation
@@
Feature Extractor None Bayar DCT FFT Sobel SRM None Bayar DCT FFT Sobel SRM
@@
Feature Extractor None Bayar DCT FFT Sobel SRM None Bayar DCT FFT Sobel SRM
@@
IMD2020 0.2679 0.0754 0.2042 0.1997 0.0844 0.2248 0.3459 0.1363 0.3223 0.2899 0.1387 0.2494
Average F1 0.466 0.143 0.364 0.364 0.147 0.363 0.538 0.214 0.472 0.447 0.225 0.395
Best Epoch 151 181 200 169 165 151 77 177 137 157 173 165
@@
Feature Extractor None Bayar DCT FFT Sobel SRM None Bayar DCT FFT Sobel SRM
CASIA v1 0.5169 0.0698 0.4001 0.4209 0.0311 0.4633 0.4195 0.2338 0.2989 0.2949 0.2031 0.4308
Columbia 0.3609 0.0854 0.2721 0.2757 0.0524 0.3057 0.1890 0.1441 0.1275 0.2656 0.1747 0.2063
NIST16 0.2344 0.1104 0.1944 0.2126 0.0861 0.1943 0.2047 0.1415 0.1560 0.1630 0.1273 0.1932
Coverage 0.1365 0.0463 0.1076 0.1097 0.0256 0.1303 0.0932 0.0585 0.0640 0.1028 0.0614 0.0824
IMD2020 0.2259 0.0284 0.1669 0.1877 0.0144 0.1755 0.1588 0.0869 0.1072 0.1149 0.0711 0.1687
Average F1 0.295 0.068 0.228 0.241 0.042 0.254 0.213 0.133 0.151 0.188 0.128 0.216
@@
A.7.3 Details of the NIST16 Dataset Cleansing Process
@@ A.7.4 Controversial F1 Metrics
Where $G$ refers to ground truth and $P$ refers to prediction. Such a formula is controversial as it does not accurately reflect the model’s detection capabilities. In Figure 4, the model’s prediction results are completely incorrect, yet its F1 score is relatively high.
@@
In the Python Skearn library, the F1 score function [^8] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.f1_score.html#sklearn.metrics.f1_score has five options for calculation methods: ‘micro’, ‘macro’, ‘samples’, ‘weighted’, ‘binary’. Except for the ‘binary’ is the commonly known F1, others are mainly used for multi-label classification problems and unsuitable for our tampering detection task with a single binary mask. This means that F1 scores are computed for each class and then averaged. Precisely, when used for binary classification, an additional F1 score is calculated by reversing both the predicted mask and the ground truth, and this score is averaged with the original F1 score. The specific averaging algorithms are different between macro and micro, and the mathematical formulas are as follows:
$F1$ $=2\times\frac{TP}{2\times TP+FP+FN}$
$F1^{\prime}$ $=2\times\frac{TN}{2\times TN+FN+FP}$
$F1_{macro}$ $=\frac{F1+F1^{\prime}}{2}=\frac{TP}{2\times TP+FP+FN}+\frac{TN}{2\times TN+FN+FP}$
$F1-F1_{macro}$ $=\frac{TP}{2\times TP+FP+FN}-\frac{TN}{2\times TN+FN+FP}$
$=\frac{1}{2+\frac{FP+FN}{TP}}-\frac{1}{2+\frac{FN+FD}{TN}}$
$F1$ $=2\times\frac{TP}{2\times TP+FP+FN}$
$F1_{micro}$ $=\frac{TP+TN}{2\times(TP+TN)+2\times(FP+FN)}$
$F1-F1_{micro}$ $=2\times\frac{TP}{2\times TP+FP+FN}-\frac{TP+TN}{2\times(TP+TN)+2\times(FP+FN)}$
$=\frac{(TP-TN)\times(FP+FN)}{(2\times TP+FP+FN)[2\times(TP+TN)+2\times(FP+FN)]}$
Since in IMDL the majority of cases typically have $TP<<TN$, it follows that $F1-F1_{macro}<0$ and $F1-F1_{micro}<0$. Consequently, $F1<F1_{macro}$ and $F1<F1_{micro}$, leading to an erroneous estimation of model performance.
We have also provided code to verify the specific implementation of macro-F1 and micro-F1 here: https://github.com/scu-zjz/IMDLBenCo/blob/main/tests/test_sklearn_F1s.py
@@ A.7.5 AUC Metrics
The Area Under the ROC Curve (AUC) is a metric used to measure the overall performance of a binary classification model across all possible classification thresholds[19]. The ROC curve itself plots the True Positive Rate (TPR) against the False Positive Rate (FPR) at various threshold settings. Visualization is shown in Figure 7.
Figure 7: Exapmple of AUC, and how trapezoid constructed when only have three data points.
Because of the threshold values, TPR and FPR are estimated through discrete points in practice. To compute the integral, we leverage the classic trapezoid rule, which involves dividing the area under the ROC curve into as many as possible trapezoids and summing their areas.
Set tuples $(FPR_{1},TPR_{1}),(FPR_{2},TPR_{2}),...,(FPR_{n},TPR_{n})$ to represent all the trapezoids
for areas under ROC, and $FPR_{i}$ is the i-th trapezoids along with a certain threshold value.
Then apply the trapezoid rule:
[…]
$\approx\sum_{i=1}^{n}\frac{TPR_{i+1}-TPR_{i}}{2}\times(FPR_{i+1}-FPR_{i})$ (3)
$=\sum_{i=1}^{n}\frac{1}{2}(\frac{TP_{i+1}}{TP_{i+1}+FN_{i+1}})\times(\frac{FP_{i+1}}{FP_{i+1}TN_{i+1}}-\frac{FP_{i}}{FP_{i}+TN_{i}})$ (4)
where i-th is the threshold for separating positive and negative samples.
We found that the AUC is calculated by accumulating the results across all thresholds. However, in practice, since the distribution of the dataset is unknown, a threshold of 0.5 is typically chosen for making predictions. This leads to integrating over unnecessary regions when the model’s ROC curve deviates significantly from 0.5, resulting in an overestimation. Such a situation is also supported by previous work (Fig.8 in MVSS-Net++[7]).
@@
Further, the manipulated regions in an image are usually very small. Therefore, in manipulated images, the number of manipulated pixels and the number of authentic pixels are extremely unbalanced. Due to such a large skew in class distribution, although AUC performs well on many unbalanced classification problems, it still tends to provide an over-optimistic estimate of IML model performance. Recent studies [4, 48] have also noticed similar issues with finding the optimal threshold, but they did not provide a profound solution. In this paper, we speculate that this disastrous performance is due to overfitting caused by large-scale pre-training, and the AUC metric is not sufficient to address this overfitting problem. Therefore, considering the overly optimistic estimates of the AUC metric in IML, we advocate using only the F1 score to evaluate IML models.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [No]
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [No]
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [Yes]
    2. (b)
      Did you include complete proofs of all theoretical results? [Yes]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes]
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes]
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No]
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes]
    2. (b)
      Did you mention the license of the assets? [No]
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [No]
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]